OpenStudy (anonymous):

x-6square root of x=16

OpenStudy (anonymous):

\[x-6\sqrt{x}=16\] \[x-16=6\sqrt{x}\] now square both sides

OpenStudy (smurfy14):

x=64

OpenStudy (anonymous):

\[(x-16)^2=36x\] \[x^2-32x+256=36x\] \[x^2-68x+256=0\]

OpenStudy (anonymous):

\[(x-4)(x-64)=0\] \[x=4\] or \[x=64\] and you have to check both answers because you squared

OpenStudy (anonymous):

4 does not work (by checking) but 64 does